OpenStudy (anonymous):

-5(3m+6)=-3(4m-2) Please hel me step by step

OpenStudy (anonymous):

Distribute

OpenStudy (anonymous):

distribute, so -5+3m = -2m -5+6 = 1 (-2m+1) -3+4m = 1m -3-2 = -5 (1m-5)

OpenStudy (anonymous):

What?!

OpenStudy (triciaal):

when you distribute to get rid of the parenthesis you multiply each term inside by what is outside the parenthesis

OpenStudy (anonymous):

So -5*3m =-15m And -5*6=-30

OpenStudy (triciaal):